OpenStudy (anonymous):

Is this correct: 2 ---x+1 = 27, which = 13 3

OpenStudy (ash2326):

you found x=13?

OpenStudy (anonymous):

Yea

OpenStudy (ash2326):

Check again, there is some mistake

OpenStudy (anonymous):

Where?

OpenStudy (ash2326):

you missed the 3 in the denominator

OpenStudy (anonymous):

Yes, you still need to times your current answer by 3 in order to get the final answer...

OpenStudy (ash2326):

@Jenovaxxxxx did you notice the mistake?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Can you show me

OpenStudy (ash2326):

certainly :)

OpenStudy (ash2326):

\[\large \frac 23 x+1=27\] \[\large \frac23 x+1-1=27-1\]\[ \large \frac23 x=26\] Multiply both sides by 3/2 \[ \large \frac23 x \times \frac 32=26\times \frac 32\] \[ \large \cancel{\frac23} x \times \cancel{\frac 32}=\cancel{26}13\times \frac 3{\cancel{2}}\] we get \[\large x=13\times 3\]

OpenStudy (ash2326):

do you understand this? @Jenovaxxxxx

OpenStudy (anonymous):

Yes

OpenStudy (ash2326):

good:)